I'm completely lost on how to solve this problem:
Find the average of these numbers:
{ 14/5, 7.38 (bar over 38), 3.69 (bar over 9) }
Is there a step by step way to solve this?

Answer by Alan3354(69443)

Find the average of these numbers:
{ 14/5, 7.38 (bar over 38), 3.69 (bar over 9) }
7.38 with the bar means it's a repeating decimal, 7.38383838...
Find the fractional value of 0.38383838... first.
n = 0.38383838...
100n = 38.38383838...
-1n = 0.38383838...
-------------------- Subtract
99n = 38
n = 38/99
--> 7 38/99
-------------
3.699999...
--> 3.7 = 3 7/10
------------------
Add 14/5 + 7 38/99 + 3 7/10
= 14/5 + 731/99 + 37/10
= (2772 + 7310 + 3663)/990
= 13745/990 = 2749/198
Divide by 3
Avg = 2749/594
=~ 4.6279